OpenStudy (anonymous):

If 34.2 grams of lithium react with excess water, how many liters of hydrogen gas can be produced at 299 Kelvin and 1.21 atmospheres? Show all of the work used to solve this problem. 2 Li (s) + 2 H2O (l) 2 LiOH (aq) + H2 (g)

OpenStudy (blurbendy):

moles Li = 34.2 g / 6.941 g/mol=4.93 the ratio between Li and H2 is 2 : 1 moles H2 = 4.93/2 = 2.47 V = nRT/p = 2.47 x 0.08206 x 299 K/ 1.21 atm = 50.0 L T = 27 + 273 = 300 K Moles CH4 = pV/RT = 1.2 x 15.8 / 0.08206 x 300=0.771 mass CH4 = 0.771 mol x 16.043 g/mol=12.4 g